### v1

# Exploring Influence Factors on LLM Suitability for No-Code Development of End User IoT Applications

Abstract.
With the increasing popularity of IoT applications, end users demand more personalized and intuitive functionality. A major obstacle for this, however, is that custom IoT functionality today still requires at least some coding skills. To address this, no-code development platforms have been proposed as a solution for empowering non-technical users to create applications. However, such platforms still require a certain level of technical expertise for structuring process steps or defining event-action relations. The advent of LLMs can further enhance no-code platforms by enabling natural language-based interaction, automating of complex tasks, and dynamic code generation. By allowing users to describe their requirements in natural language, LLMs can significantly streamline no-code development. As LLMs vary in performance, architecture, training data used, and the use cases they target, it is still unclear which models are best suited and what are the influence factors determining this fit. In particular, no-code development of IoT applications by non-technical users will have completely different demands on LLMs than, e.g., code generation for more open-ended applications or for supporting professional developers. In this paper, we explore the factors influencing the suitability of LLMs to no-code development of IoT applications. We also examine the role of input prompt language on accuracy and quality of generated applications as well as the influence of LLM training data. By conducting comprehensive experiments with a range of LLMs, we provide valuable insights for optimizing LLM-powered no-code platforms, guiding the selection of the suitable LLMs and their effective application. Our findings contribute to improving the accessibility, efficiency, and user experience of no-code IoT development, ultimately enabling broader adoption of IoT technologies among non-expert users.
Large Language Models, No-Code Development Platforms
1. Introduction
The Internet of Things (IoT) is transforming how people interact with the world by seamlessly integrating smart devices into daily life. For example, smart home applications enable users to remotely control and automate household devices, while smart healthcare applications help individuals monitor their health and fitness through wearable devices (Werner et al., 2017; Pallas et al., 2020). Personalizing IoT applications enhances usability and effectiveness by delivering smart experiences tailored to users’ unique needs and lifestyles. In smart home settings customizations include, e.g., the configurations of device behavior, automation rules and third-party integrations. For example, some users may prefer window blinds that automatically adjust based on sunlight exposure to maintain steady indoor lighting, while others might prioritize wake-up routines synchronized with natural light or security systems that adapt to their daily schedules. These personalized configurations not only enhance convenience and comfort but can also improve energy efficiency, security, and privacy (Werner et al., 2017; Bao et al., 2016).
Different end users have unique preferences, routines, and priorities, which can be difficult to capture in a one-size-fits-all IoT application. Achieving tailored functionality often requires coding knowledge, making it inaccessible to non-technical users. No-code development platforms have emerged as a solution, enabling end users to create custom applications without requiring any technical expertise (Blackstock and Lea, 2016). Although no-code platforms significantly simplify the application development process, users still need to navigate complex interfaces and workflows, understand platform-specific components and abstractions, e.g., (Bohé et al., 2021; Dantas et al., 2019; Younes et al., 2018; Batista et al., 2018), which makes widespread adoption difficult, especially for those unfamiliar with structured automation concepts. While rule-based platforms, e.g., IFTTT, reduce this barrier by providing basic automation, true interoperability across different ecosystems typically depends on API integration and programming skills. This reliance on technical skills further limits accessibility to non-technical users.
For this limitation, today’s Large Language Models (LLMs) present new opportunities for lowering the technical knowledge requirement for IoT application development (Wang et al., 2025). By leveraging natural language processing and code generation capabilities, LLMs can transform user intentions into functional codes (Wang et al., 2025), i.e., a natural language description of desired functionality is the only information required from the end users. Different LLMs vary regarding their architecture, capability, training data, and design focus. While leaderboards [^1] For example: https://artificialanalysis.ai/leaderboards/models, https://llm-stats.com/, http://huggingface.co/open-llm-leaderboard, https://huggingface.co/spaces/lmarena-ai/chatbot-arena-leaderboard compare models on general purpose tests regarding, e.g., knowledge, reasoning, or code generation, a more detailed comparison is necessary to understand the influence of varying model factors for specific tasks. For instance, no-code development of IoT applications by non-technical end users will have completely different demands on LLMs than code generation as part of a professional developer’s toolchain. To optimize the development process of no-code IoT platforms, it is critical to understand model performance differences for the task of translating natural language user input into code for IoT customization.
Specifically, in the end user IoT domain, leveraging flexible, fine-grained, and high-level abstractions from the Function-as-a-Service (FaaS) paradigm for handling code execution and operation, enables LLMs to solely focus on core functionality generation (Wang et al., 2025). Existing research has so far primarily focused on the general feasibility of integrating LLMs into no-code environments (Cai et al., 2023; Chen et al., 2024; Hagel et al., 2024; Gao et al., 2024; Wang et al., 2025), but a systematic analysis of the factors influencing the performance and suitability for different tasks is still missing.
In this paper, we systematically analyze the impact factors affecting the performance of LLM-powered no-code development for end users’ custom IoT applications. We take smart home automation as a representative IoT application domain for this and investigate the following:
  1. (1)
    The impact of the LLM choice on performance considering varying task complexity, i.e., whether certain models are better suited for IoT-specific development tasks and user interaction.
  2. (2)
    The robustness of expectable no-code platform performance for different languages, namely, Chinese and English.
  3. (3)
    The role of LLM linguistic backgrounds which the text corpora a model trained on, in shaping usability and effectiveness of the no-code development platform.
By evaluating these factors, we aim to provide insights into the design and optimization of future LLM-powered no-code IoT platforms, making IoT development more accessible and enabling non-technical users to fully harness the potential of IoT applications.
The rest of the paper is organized as follows: Section 2 gives an overview of the concepts of LLMs, no-code development platforms, and FaaS for IoT applications; it also presents related work. Section 3 describes the methodology and study design we followed while investigating the impact factors of LLM-powered no-code IoT application development and Section 4 presents the experiment results and findings. Finally, we critically discuss our findings in Section 5 before coming to a conclusion in Section 6.
2. Background and Related Work
LLMs and no-code development platforms are revolutionizing the development and deployment of IoT applications by simplifying the design process and enhancing user accessibility, while Function-as-a-Service (FaaS) further streamlines automation by enabling scalable, event-driven execution of IoT tasks.
2.1. Large Language Models
LLMs are advanced artificial intelligence models trained on vast amounts of text data to understand, process and generate human-like responses, e.g., GPT serises from OpenAI (Achiam et al., 2023), Gemini from Google (Team et al., 2023), and Llama from Meta (Dubey et al., 2024). They use deep learning techniques, particularly transformer architectures, to analyze and predict text patterns, making them highly effective for various natural language processing tasks such as translation, summarization, and providing contextually relevant responses. This makes LLMs powerful tools for understanding user’s natural language input and generating corresponding outputs for various tasks, including automation workflows in IoT applications (Ulfsnes et al., 2024; Kök et al., 2024). Their ability to translate natural language into structured logic and automate tasks makes LLMs particularly beneficial for no-code development platforms, (Chen et al., 2024; Hagel et al., 2024) enabling non-technical users to interact with systems through conversational prompts. Essentially, integrating LLMs into the no-code development platforms, the barrier to customized software development is significantly lowered, as users can generate functional code, debug issues, and optimize application logic via intuitive, language-based interfaces (Cai et al., 2023; Paliwal et al., 2024; Martins et al., 2023; Gao et al., 2024).
2.2. No-Code Development Platforms for IoT Applications
No-Code development platforms enable users to design, deploy, and modify applications through visual interfaces and pre-built drag-and-drop components, eliminating the need for coding expertise (Bohé et al., 2021; Dantas et al., 2019; Younes et al., 2018; Batista et al., 2018; El Kamouchi et al., 2023). The no-code platforms democratize software development by allowing individuals without programming knowledge to build functional applications efficiently. This approach streamlines the development process, leading to fast deployment, reducing costs, and enabling flexibility (Upadhyaya, 2023; Sufi, 2023). IoT applications encompass interconnected and diverse hardware and software components that generate and exchange data, adding complexity to the development process. No-code platforms offer advantages in the IoT domain by providing simplified integration, real-time monitoring, and automated workflows (Chen et al., 2022).
The rapid development of LLMs has opened new avenues for enhancing no-code platforms, particularly in the context of IoT. With LLM-powered no-code IoT development platforms, non-technical users can describe their desired functionalities in natural language, LLM either generate the intermediate artifacts or provide functional codes to support the application development (Martins et al., 2023; Gao et al., 2024; Wang et al., 2025).
2.3. Function-as-a-Service(FaaS) for IoT Applications
Function-as-a-Service (FaaS) is a serverless computing model that allows developers to deploy individual functions that respond to specific events or triggers, without the need to manage servers or underlying infrastructure (Bermbach et al., 2021). FaaS is highly beneficial for IoT applications due to its event-driven nature (Pfandzelter and Bermbach, 2019), i.e., events from sensors or devices can trigger automated actions, such as data processing, sending notifications, or activating other devices.
As FaaS also comes with a simple programming model, i.e., functions are typically stateless, small, and event-driven with the provider handling all operational complexity, it also significantly reduces the complexity LLMs have to handle when generating custom applications based on natural language prompts of non-technical users (Wang et al., 2025). As a result, the combination of LLMs and FaaS, thus, has the potential to disrupt the existing range of no-code platforms.
2.4. Related Work
There is research exploring the adoption of LLMs into no-code and low-code development platforms to enhance the usability, efficiency and accessibility of development.
Cai et al. (Cai et al., 2023) present an LLM-based low-code approach which incorporates multiple predefined visual programming interactions to achieve stable and controllable LLM responses. LLMs here are used to generate flowchart based on user natural language input, and users, in return, can refine the generated flowchart multiple rounds to ensure the correctness of presented workflow. Their use cases showcase the feasibility of integrating LLMs to provide solutions for text generation, solving coding tasks and providing human-like services. Gao et al. (Gao et al., 2024) propose a no-code IoT development platform by letting LLM generate trigger-action programs for IoT application, specifically, smart home application, achieving both token efficiency and result accuracy. Chen et al. (Chen et al., 2024) introduce an LLM-powered zero-code platform for template-bounded automated machine learning framework. LLM are used to enable the end-to-end model development and provide comprehensive analytical report to foster user understanding. Hagel et al. (Hagel et al., 2024) propose using LLMs to generate models from technical descriptions, improving time efficiency compared to manual modeling. The generated models are applied to the domain specific languages-based low-code platform for application customization, thus, turning the low-code platform to no-code. Their study involves technical participants working under supervision to ensure correct model generation which arises concerns about the accuracy of models generated without supervision.
These low-code and no-code approaches demonstrated the feasibility of integrating LLMs into development platforms to enhance usability and efficiency. Additionally, beyond feasibility, the underlying factors that can affect system performance are not investigated in these studies yet. This remains a critical next step question in LLMs’ suitability for no-code or low-code development, i.e., what affects the performance of LLM-powered no-code and low-code platforms.
3. Methodology
We analyze the factors influencing LLM-empowered no-code development platforms through base platform selection(Section 3.1), experimental design (Section 3.2), dataset (Section 3.3), and evaluation metrics (Section 3.4). Specifically, our experiments focus on three key aspects, i.e., model selection, variation in prompt language, and community background of LLMs.
3.1. Base No-Code Platform Selection
To assess the impact of LLMs on no-code development, we select LLM4FaaS as the base platform to perform evaluation. Specifically, LLM4FaaS leverages the high levels of abstraction of FaaS paradigm to handle code execution and operation, enabling LLMs a sole focus on core functionality generation. We reuse and extend the LLM4FaaS dataset to conduct the experiments.
3.2. Experiment Plan
To systematically evaluate the impact of LLM selection used for no-code development in IoT platforms, we conduct experiments considering multiple aspects. We select five LLMs with different strong suits in, i.e., design focus, model size, domain-specific optimization, to evaluate and compare their performance in terms of syntactic and semantic success of code production. Additionally, as the user inputs can influence the understanding of LLMs, we evaluate the model performance with two different user input languages, i.e., in Chinese and English. Finally, as LLMs trained in different linguistic environments may exhibit varying performance across languages, we explore the LLMs performance by evaluating three mainstream Chinese LLMs.
3.2.1. Model Selection
With the rapid proliferation of LLMs, it is essential to evaluate their feasibility within no-code development platforms. A thorough analysis of various LLMs can help optimize model selection strategies, ensuring these platforms achieve peak performance, efficiency, and user-friendliness. To explore this, we compare five mainstream LLMs from different aspects of consideration, i.e.,
  1. (1)
    GPT-4o, a general purpose model with advanced capacity.
  2. (2)
    GPT-4o-mini, a resource-efficient and cost-effective model, to assess if a lightweight alternative can match the performance of larger models.
  3. (3)
    Copilot, a model optimized for software development, to examine whether a domain-specialized LLM outperforms general-purpose ones.
  4. (4)
    Llama, an open-source model, to evaluate the feasibility of self-hosted or fine-tuned LLMs for customization and flexibility.
  5. (5)
    Gemini, a model with strong reasoning and multimodal capabilities, to assess its effectiveness in handling complex prompts and broader contextual understanding.
This comparison provides insights into the trade-offs between model size, design focus, and domain-specific optimization, guiding future adoption strategies for LLM-powered no-code platforms.
3.2.2. Prompt Language
The input language of LLMs can presumably have a large impact depending on the text corpora a model is trained on. Models developed in English-speaking countries, i.e., English models, will presumably be trained with more English texts while models developed in from Chinese developer teams, Chinese models, may handle Chinese user input better. To assess this language-specific impact, we conduct an experiment examining how prompt language, i.e., English and Chinese, affects LLMs performance.
3.2.3. Community Background
The consideration of prompt language choice also underscores the importance to evaluate the performance with Chinese LLMs. We select three mainstream Chinese LLMs to evaluate their performance, i.e., Alibaba Qwen, Baidu Qianfan and DeepSeek R1, which also differ in design focus.
  1. (1)
    Alibaba Qwen, a model leveraging a Mixture-of-Experts (MoE) architecture, which good at handling long-context tasks and large-scale language understanding.
  2. (2)
    Baidu Qianfan, a search-integrated LLM model optimized for Chinese natural language processing and enterprise applications.
  3. (3)
    DeepSeek R1, a reasoning-focused model, which incorporates chain-of-thought(CoT) prompting, excelling in domains, e.g., mathematics, programming, and complex multistep problem-solving.
By comparing the performance of Chinese and English LLMs, we investigate whether aligning an LLMs training corpus with the input language improves its performance of LLM-powered no-code platforms.
3.3. Dataset
Figure 1. Dataset Example: We use a LLM-powered No-Code Platform, i.e., LLM4FaaS, as the base no-code platform for evaluation. Specifically, it takes the user natural language description in combined with the project context and reference code as input to LLMs. To evaluate the accuracy of LLM-generated results, we set the ground truth based on the user requirements. The accuracy is set by comparing the output from generated results and the ground truth. Essential, the dataset consists of user requirements and the corresponding ground truth.
We use and extent an existing dataset of 26 real users natural language description for 4 smart home automation tasks in varying levels of complexity(Wang et al., 2025). Specifically, the complexity of task from simple to complex varies as follows and present an example in Figure 2:
  1. (1)
    Simple: a use case which expecting a straightforward device control function.
  2. (2)
    Medium: introduce an additional layer of complexity, where the system is expected to handle three keyword-based or time-based automation tasks.
  3. (3)
    Advanced: the scenario complexity improves by triggering smart devices based on the real-time sensor readings.
  4. (4)
    Complex: the most complex scenario, where the complexity arises from the necessity of considering device coordination, balancing user behavior with environmental conditions, and accounting for the potential diversity and uncertainty of user preferences, all simultaneously
Figure 2. User Response Example: This figure presents a real user response from the dataset, illustrating four smart home automation tasks across different complexity levels. The response showcases increasing complexity, from simple device control to intricate automation involving real-time sensor data and user-environment coordination. The original response is in Chinese, and we provide an English translation for clarity.
To enhance dataset coverage and support our experiments, we extend it in the following ways:
Input Language Variation
To assess the impact of input language on LLMs performance, we translate the original Chinese user responses into English using Google Translate [^2] https://translate.google.com/, as English is the primary language for most LLMs. This allows for a direct comparison of LLM outputs across different input languages, i.e., Chinese and English, revealing potential language biases and inconsistencies.
Ground Truth Definition
To evaluate the performance of LLM-generated results, we define the ground truth based on user responses. Specifically, we construct the ground truth based on the expected console output, derived from user responses and the intended system response. The ground truth is determined on a per-user and per-task basis, representing the expected system response for each automation task from every user. This ensures consistency in evaluation and allows for an objective comparison between LLM-generated outputs and the expected results.
Essentially, our dataset consists of user natural language descriptions of smart home automation tasks, the translated English version, and ground truth outputs serving as evaluation baseline. We show an example of dataset usage in Figure 1.
3.4. Evaluation Metrics
We first evaluate the syntactic success and semantic success rate of LLM-generated results to assess the quality of the generated code. Specifically, the syntactic success is defined as the absence of errors in the generated results, and the semantic success is defined as a complete (100%) alignment with user requirements. While these metrics can showcase the feasibility, this binary classification fails to capture varying degrees of semantic accuracy, potentially overlooking cases where the results are partially aligned with user intent.
To address this limitation, we refine the evaluation metrics that quantify semantic correctness on a continuous scale, allowing for a more granular assessment of LLM-generated outputs. In detail, the semantic accuracy rate use coverage match rate (CMR) to measure the extent to which the ground truth (GT) outputs are successfully covered by the LLM-generated outputs (MO), allowing for variations in wording and additional non-disruptive information in MO. CMR quantifies the proportion of GT entries that are correctly identified within MO. It is formally defined as:
(1) $CMR=\frac{|C|}{|GT|}$
where:
  • $|C|$ represents the number of GT entries successfully found in MO.
  • $|GT|$ denotes the total number of entries in the ground truth set.
A CMR of 100% indicates that all GT outputs are fully covered in MO, i.e., achieves semantic success, while lower scores suggest missing or incorrectly generated outputs. For example, if the light should be turned on and the curtain closed, but only one task is accomplished, this results in 50% semantic accuracy. This metric provides a more flexible assessment of model performance in real-world applications where exact textual matches may not always be required, but semantic correctness is essential.
Here, we allow the existence of additional outputs, e.g., description information before triggering the smart devices, as long as the core functionality is correctly implemented. While we acknowledge that additional outputs in LLM-generated results may introduce noise, our current evaluation focuses solely on coverage accuracy. Thus, we do not incorporate a separate noise rate metric in this study.
4. Results
We present the experimental results, analyzing the impact of key factors on the performance of LLM-powered no-code platforms. The results highlight key findings across model selection (Section 4.1), prompt language (Section 4.2), and community background (Section 4.3), providing insights into their respective impacts.
4.1. Model Selection
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 3. We consider syntactic success as error-free results and semantic success as results fully meets user requirements. The graphs depict results based on Chinese user prompts, the original language of the used dataset. GPT-4o shows distinct advantages in both syntactic and semantic success compared to other models. GPT-4o-mini performs adequately on the easy task, but the performance drops significantly on more complex ones.
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 4. GPT-4o consistently outperforms other models in semantic accuracy across tasks of varying difficulty, showing a higher accuracy rate in cases of non-semantic-success. As the task complexity increases, the semantic accuracy distribution of results becomes more dispersed. The results show a bimodal distribution, with most tasks either failing completely or achieving 100% success.
We compare the performance variance across five mainstream models, i.e., GPT-4o, GPT-4o-mini, Gemini-1.5-flash, Llama-3.1, and Copilot. First, we show the syntactic and semantic success rate of results in Figure 3, then illustrate the semantic accuracy distribution in Figure 4.
The ranking of model performances for all or most experiments remains stable in the following order: GPT-4o, GPT-4o-mini, Gemini-1.5-flash, Copilot, and Llama-3.1. Among these, GPT-4o and GPT-4o-mini show the best performance in both syntactic and semantic success. Gemini and Copilot have a similar performance for average semantic and syntactic success rate, Copilot performs better on simple tasks, and Gemini in complex tasks. Llama fails to generate functions with the user prompts. Specifically, the average syntactic success rate of GPT-4o and GPT-4o-mini is 89.10% and 73.99%, respectively, while the other LLMs are below 50%. The average semantic success rate of GPT-4o is 67.54%, whereas the second best, GPT-4o-mini, has dropped to 32.07%, and other LLMs is below 15.00%. In addition to the semantic success rate, we are also curious about how close the results are to the semantic success, given by the semantic accuracy rate. Figure 4 depicts the semantic accuracy, showing a high-level of semantic accuracy of GPT-4o which still demonstrates a clear advantage over the competing models across all difficulty levels. For all the models, the distribution of semantic accuracy of results becomes more dispersed as the task complexity increases. Notably, the results of complex tasks sometimes exhibit higher semantic success than those of medium and advanced tasks, i.e., it does not always align with the predefined task complexity. This may be due to the fact that both medium and advanced tasks involve three sub-tasks that are unrelated to each other, which can lead to more semantic failures.
4.2. Prompt Language
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 5. Success rates with English prompts. GPT-4o still outperforms other models with English prompts in both syntactic and semantic accuracy.
(a) GPT-4o
(b) GPT-4o-mini
(c) Gemini
(d) Copilot
Figure 6. Semantic success rate with prompts in Chinese and English. Translating user requirements from Chinese (ZH) to English (EN) does not significantly affect the semantic success rate. For results in both languages, OpenAI models show a clear advantage.
(a) GPT-4o
(b) GPT-4o-mini
(c) Gemini
(d) Copilot
Figure 7. Syntactic success rate with prompts in Chinese (ZH) and English (EN). The language of user prompts has little impact on the syntactic success rate. GPT-4o and GPT-4o-mini maintain a high syntactic success rate with English prompts.
To investigate the impact of prompt language to LLM performance, we use the translated English user prompt from dataset and evaluate the performance among the five LLMs. We show the syntactic and semantic success rate of English prompt results in Figure 5, then show the comparison to original results in Figure 6 and Figure 7.
GPT-4o and GPT-4o-mini, i.e., OpenAI models maintain relatively high syntactic success rates, i.e. 89.78% and 71.12%, respectively, while the other LLMs remain below 50%. Llama still fails to generate functions for all the prompts, resulting in 0% syntactic and semantic success rates.
For GPT-4o, the syntactic success rate increases slightly by 0.69 percentage points, while the semantic success rate decreases by 12.34 percentage points compared to the original prompts. English prompts improve performance only on easy tasks but negatively impact more complex ones. The average syntactic and semantic success rates of GPT-4o-mini almost remain unchanged, with a 1.68 percentage point increase in semantic success rate and a 2.87 percentage point decrease in syntactic success rate. For tasks in different complexity levels, it shows an opposite trend to GPT-4o, where it drops for easy and medium tasks but improves for advanced and complex ones. The translation step helps GPT-4o-mini to better understand the user requirements with advanced and complex tasks, while losing nuances of the original user requirements leads to a performance drop in the easy and medium tasks. Gemini experiences a decrease in both syntactic and semantic success rates by 8.77 percentage points and 6.73 percentage points, respectively. Copilot, in contrast, demonstrates an overall improvement of 9.17 percentage points in syntactic success and 5.7 percentage points in semantic success. Specifically, with English prompts, Copilot achieves a comparable performance to GPT-4o-mini with the easy task.
4.3. Community Background
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 8. Comparison of OpenAI models with Chinese LLMs: GPT-4o still outperforms other models in both syntactic and semantic accuracy, but the Chinese models show a clear advantage to GPT-4o-mini. Baidu-Qianfan can reach a similar semantic performance in easy tasks and syntactic performance with GPT-4o for all difficulty levels.
(a) Easy
(b) Medium
(c) Advanced
(d) Complex
Figure 9. Semantic Accuracy Comparison with Mandarin Models: GPT-4o also outperforms the Mandarin models in semantic accuracy across tasks of varying difficulty, but the differences are less pronounced. The two Mandarin models showcase a similar performance of semantic accuracy where Baidu-Qianfan is slightly better. With the task complexity increase, the semantic accuracy distribution of results becomes more dispersed for all models.
Considering LLMs trained with language-specific data may have a better understanding of the user prompts in the same language, we evaluate the performance of three Chinese LLMs, i.e., Baidu-Qianfan, Ali-Qwen-Max and DeepSeek-r1:7b. We compare their performance with the OpenAI models, which performed best in the initial experiments and show the results in Figure 8 and Figure 9.
Baidu-Qianfan shows a same syntactic performance to GPT-4o with an 89.10% syntactic success rate, followed by Ali-Qwen-Max, GPT-4o-mini and DeepSeek with 77.79%, 73.99% and 46.92%, respectively. In terms of the overall semantic success rate, GPT-4o maintains its lead with 67.54%, outperforming the other models. Baidu-Qianfan, Ali-Qwen-Max, and GPT-4o-mini follow with success rates of 51.93%, 43.28%, and 32.07%, respectively. DeepSeek entirely fails to generate function that meets user request, lead to a 0% semantic success rate. Notably, Baidu-Qianfan can achieve a similar semantic success rate (78.26%) to GPT-4o (82.61%) in the easy task but not for more complex ones. Ali-Qwen-Max has a comparable performance to GPT-4o-mini in the easy task but shows a clear advantage in the tasks with higher complexity. Excluding DeepSeek which has a 0% semantic success rate, GPT-4o-mini has the worst performance among this comparison, which can be caused by both the model capability and understanding of prompt language.
4.4. Findings
Feasibility of different LLMs
According to the experimental results, OpenAI LLMs, i.e., GPT-4o and GPT-4o-mini, and Chinese LLMs, i.e., Baidu Qianfan and Ali Qwen, demonstrate strong performance in both syntactic and semantic accuracy, making them promising candidates for LLM-based FaaS platforms. Specifically, GPT-4o surpasses the Chinese LLMs in semantic accuracy while achieving a comparable syntactic success rate. The Chinese LLMs, exhibit a clear advantage over GPT-4o-mini in both syntactic and semantic accuracy. Notably, GPT-4o-mini performs well on simple tasks but struggles with complex ones due to the lightweight factor. Other LLMs, i.e., Gemini, Copilot, DeepSeek, and Llama, face challenges in maintaining both syntactic and semantic accuracy.
Impact of Prompt Language
Translating user requirements from Chinese to English has only a slight impact on the syntactic performance of OpenAI models but affects semantic accuracy more strongly, leading to a performance drop for GPT-4o but an improvement for GPT-4o-mini. Although Gemini and Copilot have limited performance, translation improves both the syntactic and semantic accuracy of Copilot while reducing these of Gemini. The variance in prompts has little effect on Llama, which consistently fails to generate functions in both Chinese and English.
Alignment of LLM Performance with Task Complexity
Interestingly, the performance of the LLMs does not entirely align with the task complexity, i.e., they perform better on the complex task than the medium and advanced tasks, except for the setting of Chinese prompts used with GPT-4o. For results using translated English prompts, even GPT-4o fails to maintain the alignment. The reason for that is because both the medium and advanced tasks explicitly specify with three unrelated sub-tasks, each of which is relatively straightforward in terms of both user requirement description and the expected function logic. However, the semantic success rate drops significantly. While the complex task involves a more complex functional logic and uncertainty and diversity of user preferences, it performs better. Specifically, the Chinese LLMs exhibit similar performance alignment with what tasks specified, except for the medium task, where user requirements generally involve triggering more devices than the advanced task. These findings suggest that while the task complexity affects the LLM performance, existence of sub-tasks also matters. For LLMs weak on natural language understanding, the medium task is more challenging than the complex task. Additionally, a strong natural language understanding capability is essential for LLMs to generate correct results.
Reason for Performance Limitations
According to the experimental results, the GPT-4o and the Chinese LLMs demonstrate a strong performance in terms of syntactic and semantic accuracy. In contrast, other LLMs struggle with semantic accuracy, though the underlying causes differ.
  1. (1)
    Llama either fails to generate a function or simply restructures and rewrites the reference code provided in the prompt.
  2. (2)
    Gemini generates an excessive amount of unnecessary code snippets irrelevant to user requirements. While some generated functions contain required functionalities, they often lack semantic accuracy.
  3. (3)
    Copilot primarily generates skeletal or pseudocode structures, offering little functional implementation.
  4. (4)
    DeepSeek mainly demonstrates a thinking process of both user requirements and reference code but produces incomplete lines of code.
Overall, these results underscoring the challenges that LLMs face in generating semantically correct and functional outputs.
5. Discussion
To explore the factors influencing LLM-supported no-code development platforms, we conducted experiments using LLM4FaaS, a no-code application development platform combining the power of LLM and FaaS. Our experiments examined the impact of different model choices, the language used for prompt inputs, and the role of the LLMs’ origin communities. In this section, we discuss the implications of the findings and provide insights for future research.
5.1. Model Selection Strategy
With the release of a range of diverse LLMs in recent years, new opportunities for no-code development platforms rose. However, it is challenging to maintain an overview over model performances and select the most suitable for the use case at hand. With a vast number of available LLMs, each exhibiting different strengths and weaknesses, choosing the right model requires careful consideration. In this case, we select five LLMs with different design motivations to evaluate their performance to no-code development platforms. According to the experiment results, OpenAI models, i.e., GPT-4o and GPT-4o-mini consistently outperform the other models. This is true even for the LLM designed specific for software engineering tasks, e.g., Copilot, open-source LLM optimized for efficiency and strong adaptability, e.g., Llama, and LLM with advanced reasoning capability, e.g., Gemini. We conclude the following insights on selecting model for no-code development platforms.
5.1.1. General purpose LLM suits better than coding-specific LLMs
Copilot, as a coding-specific model should excel in generating functional code snippets. However, our experimental results indicate that Copilot does not perform well in no-code development platforms. The primary reason for its poor performance is that generating a function for non-technical users, requires more than just coding expertise, it also demands strong natural language understanding in order to interpret user prompts and project context accurately. While coding-specific models excel at code comprehension and generation, they tend to favor technical inputs over natural language descriptions from non-technical users. Our experiments show that Copilot often produces pseudocode, incomplete code skeletons, or fails to generate functional outputs altogether indicating that it struggles to transfer natural language requirements into concrete functional implementations. Additionally, Copilot frequently requires further iterations with more detailed information, highlighting its limitations in natural language understanding. In contrast, general-purpose LLMs balance both natural language understanding and code generation, making them more suitable for no-code development platforms. Our findings indicate that natural language comprehension is more critical than pure coding ability in this context, allowing general-purpose models to better interpret and implement user requirements.
5.1.2. Limitations of Lightweight LLMs for Complex Tasks
While general-purpose LLMs demonstrate superior performance in no-code development, the choice of model size and capability impacts task effectiveness. For simpler tasks, even a lightweight model, i.e., GPT-4o-mini performs well, efficiently handling straightforward logic and basic automation. Its lower computational cost makes it a viable option for quick iterations and low-complexity scenarios. However, as task complexity increases, even when the generated function remains relatively short, the lightweight models struggle to maintain accuracy and coherence. Complex workflows and nuanced logic require deeper contextual understanding and more robust reasoning abilities where requires a more powerful model to consistently deliver accurate and reliable results.
5.1.3. Performance Variations Among General-Purpose LLMs
Our experiment evaluates different general-purpose LLMs, i.e., OpenAI models, Gemini, and Llama, using the same structured prompt. The results reveal significant differences in their responses and only OpenAI models prove feasible for no-code development platforms. The other models generate lengthy but invaluable outputs. Gemini tends to produce excessive additional code snippets beyond the requested function. While this broader output can be useful, for our experiment, it introduces unnecessary complexity, incurs unintended errors or warnings, and leads to higher token consumption. Rather than improving the final code generation, the additional content requires users to manually extract relevant portions, which demands technical expertise and makes it less suitable for no-code development platforms. This behavior suggests that Gemini prioritizes completeness over precision. Llama, on the other hand, places focus on analyzing project context in the prompt. Instead of directly generating the requested function, it tends to rewrite or restructure existing reference code, often deviating from the explicit intent of the prompt. This indicates that Llama prioritizes context-aware code generation, which may be beneficial in some software engineering workflows but is less aligned with the needs of no-code development.
5.1.4. Community background of LLMs matters
As our end user descriptions origin in Chinese, we explore performance of comparable Chinese LLMs from Alibaba, Baidu, and DeepSeek. The Alibaba Qwen and Baidu Qianfan models demonstrated similar performance to the OpenAI models, while DeepSeek-r1-7B performed poorly in comparison. In detail, Qianfan slightly outperformed Qwen, and both surpassed GPT-4o-mini. While GPT-4o outperformed the Chinese LLMs in terms of semantic success, the two Chinese models demonstrated comparable performance to GPT-4o in terms of semantic accuracy, emphasizing the influence of community alignment to the result accuracy. The DeepSeek model generated results similar to those of Gemini and Llama, offering a comprehensive thinking process rather than directly producing the requested function. It can be useful for developers to understand code or for educational purposes, but it is not ideal for no-code development platforms, where more targeted and functional code generation is required. The results imply that even the linguistic background of LLMs influences their performance in no-code development platforms, the design focus of the model and precise output are the key determinant of success.
5.1.5. Response time varies among LLMs
The runtime of query completion is another important factor in the choice of model. Our experiments indicated large differences between the models, however, we did not perform specific measurements on the computation time. This would be an interesting aspect for future research.
5.2. Prompt Construction Concerns
The language used in the prompt input plays a crucial role in determining the LLM’s performance in no-code development platforms. We evaluate the performance of LLM-powered no-code platform with original Chinese prompts and translated English prompts. We find that achieving optimal performance requires adapting prompt design when applying different LLMs to no-code development platforms, as each model benefits from a distinct approach to maximize performance and ensure accurate code generation. For a general-purpose model with strong natural language understanding, e.g., GPT-4o, the original user prompt tends to work best, as the model can capture nuanced details from the description. For a lightweight model, e.g., GPT-4o-mini, the optimal strategy is the opposite. For a model with advanced reasoning capabilities that generates excessive results, such as Gemini, the original user prompt is preferred as it retains more detailed information, enabling the generation of comprehensive results. Retaining the original prompt is more effective for easy tasks, while translating the prompt into English benefits complex tasks by allowing the model to focus more on code generation rather than natural language interpretation. For a coding-specific model such as Copilot, English prompts generally yield better results than original Chinese ones. When provided with Chinese input, we notice that Copilot first translates the text into English before generating code, which can dispread the focus and introduce inaccuracies. However, with English prompts, Copilot is more likely to generate pseudocode or incomplete functions, limiting its suitability for no-code development. Although different LLMs respond differently to prompt language, the overall impact on performance remains limited. Notably, GPT-4o consistently outperforms other models, making it the most reliable choice for no-code development. These findings indicate that while prompt language can influence model behavior, strong natural language understanding is the key determinant of success.
5.3. Implications for Other Research Fields
This work focuses on understanding the performance of LLMs in generating IoT automation software for non-technical users. Beyond the question of how well LLMs can perform these tasks and what causes their behavior, introducing LLM-based automation for the IoT raises interesting research questions outside the domain of computer systems research.
For example, our proposal will require additional research in human-computer interaction. The usability and accessibility of LLMs in the domain of IoT automation should be investigated further, focusing not just on the performance of LLMs but also investigating the perception that users have of the system. It is an open question whether an unsuccessful request leads to frustration with the system or more interest and interaction as users are given an opportunity to refine their prompts. Similarly, the cognitive load on users as they formulate their intentions as well as their trust in the system could be quantified.
As we advocate for empowering more users to build their own smart home automations (through LLMs rather than writing code), we believe that there could also be effects on behavior on a broader scale. For individual users, being able to leverage the IoT even without a technical background could lead to both an increased reliance on LLM-based coding (without a desire to learn more about the underlying mechanisms) or more interest in technology as the user has positive interactions with the system. Here, it could also be interesting to evaluate what mental models users have of the system and how it influences their behavior and digital literacy, especially when taking into account social and cultural factors.
Finally, we believe that this research area also warrants ethical considerations. Leveraging LLMs to automate the smart home can have privacy implications as data on user behavior and homes is shared with the system and, potentially, an LLM service. Then, there are concerns around responsibility when executing LLM-generated code, especially in a sensitive setting such as the private home. If an IoT automation task could harm a user or third-party, e.g., by ignoring air quality sensor data or setting the heating to a dangerous level, and the user who prompted the automation does not have sufficient technical expertise to validate the LLM output, there is the question of accountability. Lastly, there is the need for further research into bias and fairness. A LLM may exhibit particularly bad behavior when confronted with automation tasks that do not fit its model of the world, e.g., when building automation around certain cultural practices.
6. Conclusion
LLM-supported no-code platforms leverage the natural language understanding capabilities of LLMs to generate functional code based on user inputs, enabling seamless customization without requiring any technical expertise. Understanding the factors influencing the performance of LLM-powered no-code platforms can give insights on how to optimize the real-world development process, ensuring reliability, optimizing efficiency and enhancing the user experience.
In this paper, we aim to find the influencing factors that affect the performance of LLM-powered no-code IoT platforms, considering of LLM choice, prompt language variance, and LLM linguistic background. We provide valuable insights into the design and optimization of future platforms. Specifically, model selection has the most significant impact on performance. A general-purpose LLM with advanced natural language understanding, which prioritizes providing concrete results over comprehensive ones, is generally preferred. A LLM which the linguistic background aligns with input language also showcase an outperforming performance. Additionally, the influence of prompt language varies across different LLMs and task complexity levels, which should be carefully considered in the design of no-code platforms. Specifically, for LLM with advanced natural language understanding and multilingual capabilities, e.g., GPT-4o, the original user prompts should be kept to ensure the best performance. While for LLMs with either in lightweight or with limit in natural language understanding capability, i.e., GPT-4o-mini and Copilot, adding a translation step can improve the performance.
Acknowledgements.
Partially funded by the Sponsor Bundesministerium für Bildung und Forschung (BMBF, German Federal Ministry of Education and Research) https://www.bmbf.de/bmbf/en in the scope of the Software Campus 3.0 (Technische Universität Berlin) program – Grant #01IS23068.